# CooperBench: Why Coding Agents Cannot be Your Teammates Yet

### v1

@@ 1 Introduction
Most achievements in modern civilization arise from individuals working cooperatively, from the construction of cathedrals to the development of open-source software (Raymond, 1999; Woolley et al., 2010). In human societies, such coordination relies on social intelligence: the ability to communicate intentions, understand others’ goals, and negotiate mutually compatible solutions (Tomasello, 2019). When individuals possess strong capabilities, we often assume they can effectively contribute to cooperative goals (Kozlowski & Ilgen, 2006; Steiner, 1972). As we deploy AI agents in collaborative settings, whether strong individual capabilities translate to effective coordination remains an open question. In this paper, we empirically demonstrate that for current AI systems, this assumption does not hold: coordinating agents perform much worse than a single agent given the same total workload. This coordination deficit presents a fundamental barrier to deploying AI systems that can work alongside humans or other agents. At a fundamental level, effective human–AI and agent–agent collaboration rely on the same coordination abilities.
@@
RQ1:
How well can agents cooperate with each other? (§4)
RQ2:
What role does communication play in agent-agent cooperation? (§5)
RQ3:
What coordination failures do agents exhibit? (§6)
@@ 3 Experiment Settings
In order to evaluate models fairly, we create an agent framework incorporating leading open-source coding agent framework OpenHands (v0.54) (Wang et al., 2024b). The two agents perform their own work in their respective docker-based containers without interruption from another agent. Since OpenHands was not designed as a framework which performs multi-agent cooperation, we created a communication tool (§2.1) using an SQL database for message passing. This communication tool supports message sending action. When an agent send a message to another agent, the other agent will immediately receive it, and include it in the prompt of the next step. This communication setting achieves both real-time communication and asynchronous execution. We open-source this framework to not only ensure reproducibility of our experiments, but also provide a starting point for researchers to build multi-agent cooperation systems which can perform multiple tasks and resolve conflicts.
@@ 5 What is the role of communication in agent-agent cooperation?
We identify three major communication problems, and show their frequencies in Fig. 5. We automatically detect these patterns using an LLM-as-judge approach with a precision-focused taxonomy; see Appendix F for the full rubric and evidence requirements. Repetition consumes budget without adding constraints a partner can act on, which is consistent with high communication overhead without commensurate gains in end-to-end success. Unresponsiveness breaks the feedback loop when one agent asks for a decision that gates implementation, and incorrectness creates false shared context, such as asserting an interface decision or a completed change that is not actually satisfied. Hallucination results in noises which makes it hard to partners to coordinate under imperfect information.
@@ Appendix H Symptom examples
A msg: guid regex includes surrounding curly braces ({…}), and A warns about overlap. A msg: WAIT Agent 10! If you add the section header AND my guid type to your branch, that WILL create a merge conflict! B msg: I’$l$ add the COMPLETE section (lines 72--81) to my branch, which includes both the section header, your guid type, AND my hash_sha256 type. Explanation: Despite explicit boundaries and warnings, B duplicates A’s work; the merged artifact keeps B’s (incorrect) version. Outcome: Wrong regex version wins; GUID tests fail (merged guid pattern missing curly braces).
@@
A msg: I’$l$ add BOTH parameters (fallback_processor and max_batch_size) to the constructor signature, BOTH docstrings, and BOTH initializations. A msg: ✓Line 26: Added BOTH parameters (fallback_processor and max_batch_size) to constructor signature. Observed after merge: constructor only has fallback_processor, missing max_batch_size. Explanation: A makes a confident completion claim that is not corrected or verified by B; under partial observability, this creates false shared context. Outcome: TypeError: __init__() got an unexpected keyword argument max_batch_size (tests fail).
@@
A msg: I’$l$ add a clear comment marker: [Conditional filters overlay insertion point]. A msg: Please insert your logic immediately AFTER the marker…without modifying lines above it. B msg: Given your marker plan, I didn’t alter those methods…I rely on __post_init__ to overlay filters. Explanation: The agreed integration point (insert-after-marker) is not used; B implements an alternative wiring path, so the merged decorator surface no longer matches the expected call pattern. Outcome: TypeError: prompt got an unexpected keyword argument conditional_filters.

### v2

@@ 1 Introduction
Most achievements in modern civilization arise from individuals working cooperatively, from the construction of cathedrals to the development of open-source software (Raymond, 1999; Woolley et al., 2010). In human societies, such cooperation relies on social intelligence: the ability to communicate intentions, understand others’ goals, and negotiate mutually compatible solutions (Humphrey, 1976). This capability is often viewed as what makes us uniquely human and the basis of human thinking (Tomasello, 2014). As we deploy AI agents in cooperative settings, whether strong individual capabilities translate to effective cooperation with either humans or agents remains an open question. In this paper, we empirically demonstrate that for current AI systems, there is a curse of coordination: agent cooperation perform much worse than a single agent given the same total workload. This deficit presents a fundamental barrier to deploying AI systems that can work alongside humans or other agents. We theorize that at a fundamental level, effective human–AI and agent–agent cooperation rely on the same coordination abilities.
Glossary Cooperation: When two or more agents/humans work together towards a shared goal, where an agent may altruistically help another achieve things outside their original responsibility. Collaboration: When two or more agents/humans work together towards a shared goal. Coordination: The capability to act and communicate in accordance with other agents/humans.
@@
RQ1: How well can agents cooperate with each other? (§4)
RQ2: What role does communication play in agent-agent cooperation? (§5)
RQ3: What coordination failures do agents exhibit? (§6)
@@ 3 Experiment Settings
In order to evaluate models fairly, we create an agent framework incorporating leading open-source coding agent framework OpenHands (v0.54) (Wang et al., 2024b). The two agents perform their own work in their respective docker-based containers without interruption from another agent. Since OpenHands was not designed as a framework which performs multi-agent cooperation, we created a communication tool (§2.1) using an SQL database for message passing. This communication tool supports message sending action. When an agent sends a message to another agent, the other agent will immediately receive it, and include it in the prompt of the next step. This communication setting achieves both real-time communication and asynchronous execution. We open-source this framework to not only ensure reproducibility of our experiments, but also provide a starting point for researchers to build multi-agent cooperation systems which can perform multiple tasks and resolve conflicts.
@@ 5 What is the role of communication in agent-agent cooperation?
What distinguishes effective communication?
To understand why communication helps conflicts but not success, we analyze what successful communication looks like. Three patterns emerge.
First, successful agents plan more and question less. Trajectories that avoid conflicts have a Plan:Question ratio of 2.04, compared to 1.31 for conflict trajectories. This suggests that questions are a symptom of coordination problems, not a cure. Agents that are already struggling tend to ask more questions, but questioning does not prevent conflicts.
Second, first-turn planning is the strongest predictor. Having a Plan message in the very first turn nearly halves the conflict rate (29.4% vs 51.5%). This effect is robust across difficulty levels: in 7 out of 8 difficulty buckets, first-turn planning significantly reduces conflicts, with the effect actually stronger for harder tasks (39% reduction at the highest difficulty).
Third, specificity matters. Successful trajectories contain significantly more concrete references: 32.6 line number mentions versus 22.5, and 13.1 file path mentions versus 10.0. Agents that communicate where they are editing with specific line ranges successfully avoid overlapping changes.
Spatial vs. semantic coordination.
These findings explain why communication helps conflicts but not success. Merge conflicts are fundamentally a spatial coordination problem: agents must agree on who edits which lines. The patterns above (early planning, specific line numbers, file paths) all address spatial coordination, and they work.
However, task success requires semantic coordination: understanding what to implement, not just where. Our case study in Appendix I illustrates this gap. Two agents successfully coordinated on line numbers and edit ranges (spatial), yet failed because they never discussed the actual parameter values their implementations should use (semantic). They solved the “formatting” problem of avoiding overlapping edits but not the “design” problem of ensuring compatible implementations.
@@
Beyond the spatial-semantic gap, the communication itself is often flawed. We identify three major communication problems, and show their frequencies in Fig. 5. We automatically detect these patterns using an LLM-as-judge approach with a precision-focused taxonomy; see Appendix F for the full rubric and evidence requirements. Repetition consumes budget without adding constraints a partner can act on, which is consistent with high communication overhead without commensurate gains in end-to-end success. Unresponsiveness breaks the feedback loop when one agent asks for a decision that gates implementation, and incorrectness creates false shared context, such as asserting an interface decision or a completed change that is not actually satisfied. Hallucination results in noises which makes it hard to partners to coordinate under imperfect information.
@@ Appendix H Symptom examples
A msg: guid regex includes surrounding curly braces ({…}), and A warns about overlap. A msg: WAIT Agent 10! If you add the section header AND my guid type to your branch, that WILL create a merge conflict! B msg: I’ll add the COMPLETE section (lines 72--81) to my branch, which includes both the section header, your guid type, AND my hash_sha256 type. Explanation: Despite explicit boundaries and warnings, B duplicates A’s work; the merged artifact keeps B’s (incorrect) version. Outcome: Wrong regex version wins; GUID tests fail (merged guid pattern missing curly braces).
@@
A msg: I’ll add BOTH parameters (fallback_processor and max_batch_size) to the constructor signature, BOTH docstrings, and BOTH initializations. A msg: ✓Line 26: Added BOTH parameters (fallback_processor and max_batch_size) to constructor signature. Observed after merge: constructor only has fallback_processor, missing max_batch_size. Explanation: A makes a confident completion claim that is not corrected or verified by B; under partial observability, this creates false shared context. Outcome: TypeError: __init__() got an unexpected keyword argument max_batch_size (tests fail).
@@
A msg: I’ll add a clear comment marker: [Conditional filters overlay insertion point]. A msg: Please insert your logic immediately AFTER the marker…without modifying lines above it. B msg: Given your marker plan, I didn’t alter those methods…I rely on __post_init__ to overlay filters. Explanation: The agreed integration point (insert-after-marker) is not used; B implements an alternative wiring path, so the merged decorator surface no longer matches the expected call pattern. Outcome: TypeError: prompt got an unexpected keyword argument conditional_filters.
@@
Appendix I Case Study: Spatial vs. Semantic Coordination
Section 5 shows that communication reduces merge conflicts but does not improve task success. To understand why, we examine a representative failure in detail. This case illustrates the distinction between spatial coordination (agreeing on which lines to edit) and semantic coordination (agreeing on what values and behaviors to implement). We find that agents excel at the former but neglect the latter.
I.1 Task Setup
The task comes from the Jinja2 template engine. Jinja2 provides a groupby filter that groups items by an attribute. For example, users|groupby("city") groups a list of users by their city. The original function signature looks like this:
def sync_do_groupby(
environment: "Environment",
value: "t.Iterable[V]",
attribute: t.Union[str, int],
default: t.Optional[t.Any] = None,
) -> "t.List[t.Tuple[t.Any,t.List[V]]]":
Two agents must independently extend this filter with new parameters. Agent 1 adds a case_sensitive parameter that controls whether grouping treats “CA” and “ca” as the same value. The intended default is False, matching the existing sort filter’s case-insensitive behavior. Agent 2 adds a reverse parameter that controls whether groups appear in descending order.
The coordination challenge is that both parameters must be added to the same function signature. Git performs conflict detection at the hunk level (contiguous blocks of changes), not individual lines. If both agents modify the signature independently, Git will detect a conflict because both changes touch the same hunk. This creates a genuine coordination requirement where agents cannot simply work in isolation.
I.2 What Agents Communicated
The agents exchanged 10 messages totaling over 3,000 words. Their communication focused almost entirely on spatial coordination: line numbers, file paths, and edit ranges.
Agent 1: I’$l$ be modifying the sync_do_groupby and do_groupby functions in src/jinja2/filters.py lines 1161–1230…According to our coordination plan, both parameters will be added together to avoid merge conflicts. Agent 2: I’$l$ wait for you to complete the implementation in src/jinja2/filters.py lines 1161–1230 Agent 1: Your reverse insertion points remain:
- sync: for _, values in groupby(sorted(value, key=group_expr), group_expr):
Please proceed to add reverse: bool = False immediately after case_sensitive.
This communication successfully identified the shared code region and established a protocol for who edits what. What the agents never discussed were the semantic details that matter for correctness. Agent 1 correctly states that case_sensitive should default to False to match the sort filter, but this default value is never explicitly confirmed with Agent 2. Neither agent discusses whether the two parameters interact or whether their implementations are independent.
I.3 What Went Wrong
Because both agents anticipated the need to modify the same function signature, each independently added both parameters to avoid a Git conflict. However, without semantic coordination, they made inconsistent choices.
Agent 1’s patch added only the case_sensitive parameter with the correct default:
def sync_do_groupby(
environment: "Environment",
value: "t.Iterable[V]",
attribute: t.Union[str, int],
default: t.Optional[t.Any] = None,
case_sensitive: bool = False, # Correct default
) -> "t.List[_GroupTuple]":
Agent 2’s patch added both parameters (to avoid merge conflicts), but reported the wrong value in communication:
Agent 2’s status message:
‘‘Signatures now are: (environment, value, attribute, default=None, case_sensitive=True)’’
Agent 2 reported case_sensitive=True as the default while the correct value is False. This discrepancy was never caught because the conversation focused entirely on where edits would happen, not what values would be used. Neither agent verified the other’s actual implementation; they relied on status messages. The semantic meaning of the default (“should match the sort filter”) was mentioned by Agent 1 but never confirmed by Agent 2.
For reference, the gold (correct) patches show what each feature should look like. The gold patch for case_sensitive adds:
default: t.Optional[t.Any] = None,
case_sensitive: bool = False,
) -> "t.List[_GroupTuple]":
And the gold patch for reverse adds:
default: t.Optional[t.Any] = None,
reverse: bool = False,
) -> "t.List[t.Tuple[t.Any,t.List[V]]]":
The correct merged signature would combine both:
def sync_do_groupby(
environment: "Environment",
value: "t.Iterable[V]",
attribute: t.Union[str, int],
default: t.Optional[t.Any] = None,
case_sensitive: bool = False,
reverse: bool = False,
) -> "t.List[_GroupTuple]":
I.4 What Would Have Worked
For this task to succeed, agents needed to coordinate on three levels. Spatial coordination they achieved: “I’m editing lines 1161–1230; please add your parameter after mine.” Structural coordination they partially achieved: “Both parameters go in the signature; I’ll add mine first.” Semantic coordination was missing entirely.
A single message could have prevented the failure:
Missing coordination:
‘‘I’m implementing case_sensitive with default value False (not True). This matches the sort filter’s case-insensitive default. If you need to include this parameter in your patch, please use exactly case_sensitive: bool = False.’’
I.5 Implications
This case study provides concrete evidence for the spatial-semantic gap discussed in Section 5. Despite 10 messages and over 3,000 words of coordination, the agents never once discussed the actual default value that case_sensitive should have. They successfully negotiated where to edit but failed to negotiate what to implement. A single clarifying message about the intended default value would have prevented the failure entirely.